What power is needed to lift a 49-kg person a vertical distance of 5.0 m in 20.0 s?
1
Expert's answer
2013-11-26T06:24:56-0500
Solution:
Power=Work/time
Work=Force*Distance (in our case)
Force=m*g (lifting force is equal to weight of the body)

now numbers
m=49kg
Distance=5m
time=20s
g=10H/kg

Force=m*g=49kg*10N/kg=490N
Work=Force*distance=490N*5m=2450Nm
Power=Work/time=2450Nm/20s=122.5Nm/s=122.5Watt

ANSWER:
Power=122.5Watt